# Enhancing the Reliability in Machine Learning for Gravitational Wave Parameter Estimation with Attention-Based Models

### v1

@@ Abstract
We introduce a technique to enhance the reliability of gravitational wave parameter estimation results produced by machine learning. We develop two independent machine learning models based on the Vision Transformer to estimate effective spin and chirp mass from spectrograms of gravitational wave signals from binary black hole mergers. To enhance the reliability of these models, we utilize attention maps to visualize the areas our models focus on when making predictions. This approach enables demonstrating that both models perform parameter estimation based on physically meaningful information. Furthermore, by leveraging these attention maps, we demonstrate a method to quantify the impact of glitches on parameter estimation. We show that as the models focus more on glitches, the parameter estimation results become more strongly biased. This suggests that attention maps could potentially be used to distinguish between cases where the results produced by the machine learning model are reliable and cases where they are not.
@@ I Introduction
In [18] and [19], Bayes’ rule is applied to enable network training using only samples from the prior distribution and gravitational wave likelihood. This approach allows for the estimation of posterior distributions without the need for posterior samples during training. As noted above, likelihood evaluation requires significant computational costs. Therefore, these so-called “likelihood-free“ methods are highly valuable, leading to the exploration of many potential applications [20, 21, 22]. A notable example is the DINGO framework [23], which uses normalizing flows to estimate all 15 parameters of BBH mergers, achieving results consistent with the posterior distributions obtained from LALInference. In the DINGO framework, the neural network is conditioned not only on the strain data but also on the power spectral density (PSD) of the detector noise, effectively addressing the fact that detector noise varies from event to event. This excellent framework has also been expanded to parameter estimation for binary neutron stars [24]. Therefore, machine learning is progressing toward practical applications not only in gravitational wave detection and glitch classification but also in parameter estimation (See, e.g., [25, 26, 27, 28]). (For typical approaches that do not use machine learning, such as reduced-order modeling, efficient sampling methods, and others, see, e.g., [29, 30, 31, 32] and references therein.)
When applying machine learning to gravitational wave parameter estimation, it is important to consider the “black-box“ problem [33, 34]. Because machine learning models learn from large amounts of data and autonomously generate outputs, their decision-making process is unclear. For example, even if there are significant noises such as glitches nearby gravitational wave signals, machine learning models generate outputs without warning. For another example, it has not been sufficiently verified whether the models are focusing on physically meaningful information.
In this work, we visualized attention maps of a Vision Transformer(ViT) [35]. Attention maps show the areas a machine learning model focuses on when it makes predictions, thereby improving the reliability of outputs.
@@ II.1 GW dataset generation
For the generation of CBC waveforms, we used PyCBC [36] with the SEOBNRv4_opt [37] waveform model. Our models were trained with parameters from Tab 1. Although some parameters are restricted, we chose a simple configuration for the sake of simplicity, considering our main purpose.
Regarding the data generation procedure, we first created strain data by injecting a waveform into 4 second long data of stationary Gaussian noise colored by the LIGO design sensitivity (aLIGOZeroDetHighPower) [36]. During this process, the central time of the noise data was aligned with the merger time of the waveform. Then, we whitened the strain and applied a constant-Q transform to create the spectrogram. (In the whitening process, we cut the first and last 0.5 seconds of data to minimize possible edge effects.)
@@
Table 1: Parameters used to create the training and validation datasets. $\mathcal{M}$, $q$, $\chi_{\text{eff}}$, $\theta_{JN}$, and $\psi$ are uniformly distributed. Source-frame component masses $m_{1}$,$m_{2}$, are determined from $\mathcal{M}$ and $q$, while the dimensionless projections of the individual BH spins $\chi_{1}$ and $\chi_{2}$ are determined from $\chi_{\text{eff}}$. The luminosity distance $d_{L}$ and the sky position (right ascension $\alpha$ and declination $\delta$) are set to ensure a uniform distribution of gravitational wave sources within a sphere.
@@ II.2 Training and validation
We used Data-efficient Image Transformers (DeiT) [38], which is a model based on ViT. DeiT successfully reduces the computational cost of training while maintaining accuracy by utilizing knowledge distillation [39] and data augmentation. We used DeiT models pre-trained on ImageNet [40] and trained it on the dataset we created.
@@ II.3 Evaluation of uncertainty
To evaluate the uncertainty in the predictions of our models, we developed an uncertainty evaluation method based on a primitive Monte Carlo approach. The flowchart of this method is shown in Fig. 3.
(1) Initially, a waveform was generated using parameters consistent with the estimates obtained from LALInference for GW150914. After injecting this waveform into stationary Gaussian noises colored by the same PSD as in section II.1, a constant-Q transform was applied to create a spectrogram.
@@
(3) Subsequently, a waveform was generated based on the values estimated in (2). For parameters that were not estimated, the values were set to those used in (1). This waveform was injected into stationary Gaussian noises colored by the same PSD as in section II.1, and a constant-Q transform was applied to create spectrograms. In this way, we created an uncertainty evaluation dataset that contains 1,000 spectrograms.
@@ III.1 Prediction results
Tab. 2 shows the 90% confidence intervals, while Fig. 4 and 5 illustrate the distributions of absolute errors. In Tab. 2, it can be seen that the predictions of our models were approximately consistent with those obtained by LALInference for GW150914. However, we did not analyze real GW events, and since the PSD used corresponds to the design sensitivity, this represents the result of a rough comparison.
@@
Figure 4: Histogram of absolute error for the effective spin estimation model. The absolute errors are calculated from the effective spins estimated in section  II.3-(4) and the estimated value in section  II.3-(2).
Figure 5: Histogram of absolute error for the chirp mass spin estimation model. The absolute errors are calculated from the chirp masses estimated in section  II.3-(4) and the estimated value in section  II.3-(2).
@@ III.2 Attention Map
We generated attention maps for the 1,000 spectrograms used in the uncertainty evaluation described in section  II.3. To illustrate the overall trend, we created a composite figure by overlaying these attention maps (Fig. 6). It can be seen that the effective spin estimation model appears to focus primarily on the mid-frequency range, whereas the chirp mass estimation model focuses on the low-frequency range.
@@ IV.1 Method
We injected glitches into the simulated GW150914 data and applied the same uncertainty evaluation method as in section  II.3 to estimate the extent to which our model’s predictions would be biased if glitches were present in the simulated GW event. Furthermore, we examined the relationship between the Attention Map and the prediction results.
@@
$h_{\rm sg}=A_{\rm sg}e^{-\frac{(t-t_{0})^{2}}{Q}}\cos\left(2\pi f_{0}(t-t_{0})\right)$ (1)
$A_{\rm sg}$ is an overall amplitude, $t_{0}$ and $f_{0}$ are the location of the sine-Gaussian in time and frequency, respectively, Q is the dimensionless quality factor.
In real GW data analysis, it is problematic when glitches directly overlap GW signals as in the case of GW170817 [15]. However, injecting around the merger phase seemed likely to cause significant bias, thereby making it difficult to identify clear trends. Therefore, we selected $Q=10^{-4}$, $f_{0}=40$ [Hz] and $t_{0}=1.0$ [s] to configure the glitches to slightly overlap with the inspiral phase. For the amplitude $A_{\rm sg}$, we prepared three glitches with amplitudes of $0.5A$, $0.8A$, and $1A$, using the maximum amplitude $A$ of the simulated GW signal. In this manner, we generated spectrograms for three simulated GW events with injected glitches. Fig. 8 presents an example.
@@ IV.2 Predictions
We compared the prediction results from section  III.1 with three simulated GW events with glitches of three different amplitudes.
Tab. 3 shows the 90% confidence intervals, while Fig. 9 and 10 illustrates the distributions of the predicted values. It can be seen that for both effective spin and chirp mass, glitches with larger amplitudes more strongly biased parameter estimation. This tendency is also commonly seen in conventional methods.
@@
$\chi_{\text{eff}}$ $-0.09^{+0.07}_{-0.09}$ $-0.09^{+0.09}_{-0.07}$ $-0.07^{+0.08}_{-0.07}$ $0.04^{+0.07}_{-0.05}$
$\mathcal{M}/M_{\odot}$ $27.6^{+0.8}_{-1.0}$ $30.5^{+0.8}_{-1.1}$ $32.6^{+0.6}_{-1.4}$ $33.3^{+0.7}_{-1.3}$
@@ IV.3 Attention Map
Fig. 11 and 12 shows the Attention Maps for a simulated GW150914 event without glitches and for three simulated GW events with varying glitch amplitudes. It can be seen that as the amplitude of the glitch increases, the attention around 1.0 seconds and 40 Hz, where the glitch was injected, also becomes larger.
@@
Furthermore, to quantify the impact of glitches to some extent, we calculated the total attention values for each time point. The distribution is shown in Fig. 13 and 14. The figures illustrate that as the amplitude of the glitch increases, the attention values around it also tend to increase.
@@ V Conclusion
VI Acknowledgements
We would like to thank Nobuyuki Kanda, Masaki Uematsu, Izumi Kaku, Takumi Fujimori, Ryuki Kawamoto, and Hiroyuki Nakano for helpful comments. We are grateful to Kota Tomita and Kyohei Takatani for their contribution in the early phase of this work. We would also like to thank Hirotaka Takahashi, Takashi Uchiyama, Yutaka Shikano, Shoichi Oshino and Yusuke Sakai for the valuable insights provided regarding the application of machine learning to gravitational wave data analysis. This work was supported by JSPS KAKENHI Grant Number 20H05639. We use PyTorch [56] for the implementation of our models. The plots are generated with matplotlib [57].

### v2

# Enhancing the reliability of machine learning for gravitational wave parameter estimation with attention-based models

@@ Abstract
We introduce a technique to enhance the reliability of gravitational wave parameter estimation results produced by machine learning. We develop two independent machine learning models based on the Vision Transformer to estimate effective spin and chirp mass from spectrograms of gravitational wave signals from binary black hole mergers. To enhance the reliability of these models, we utilize attention maps to visualize the areas our models focus on when making predictions. This approach enables us to demonstrate that both models perform parameter estimation based on physically meaningful information. Furthermore, by leveraging these attention maps, we demonstrate a method to quantify the impact of glitches on parameter estimation. We show that as the models focus more on glitches, the parameter estimation results become more strongly biased. This suggests that attention maps could potentially be used to distinguish between cases where the results produced by the machine learning model are reliable and cases where they are not.
@@ I Introduction
In [18] and [19], Bayes’ rule is applied to enable network training using only samples from the prior distribution and gravitational wave likelihood. This approach allows for the estimation of posterior distributions without the need for posterior samples during training. As noted above, likelihood evaluation requires significant computational costs. Therefore, these so-called “likelihood-free” methods are highly valuable, leading to the exploration of many potential applications [20, 21, 22]. A notable example is the DINGO framework [23], which uses normalizing flows to estimate all 15 parameters of BBH mergers, achieving results consistent with the posterior distributions obtained from LALInference. In the DINGO framework, the neural network is conditioned not only on the strain data but also on the power spectral density (PSD) of the detector noise, effectively addressing the fact that detector noise varies from event to event. This excellent framework has also been expanded to parameter estimation for binary neutron stars [24]. Therefore, machine learning is progressing toward practical applications not only in gravitational wave detection and glitch classification but also in parameter estimation (See, e.g., [25, 26, 27, 28]). (For typical approaches that do not use machine learning, such as reduced-order modeling, efficient sampling methods, and others, see, e.g., [29, 30, 31, 32] and references therein.)
When applying machine learning to gravitational wave parameter estimation, it is important to consider the “black-box” problem [33, 34]. Because machine learning models learn from large amounts of data and autonomously generate outputs, their decision-making process is unclear. For example, even if there are significant noises such as glitches nearby gravitational wave signals, machine learning models generate outputs without warning. For another example, it has not been sufficiently verified whether the models are focusing on physically meaningful information.
In this work, we used Data-efficient Image Transformers (DeiT) [35], which is a model based on ViT [36]. ViT applies the transformer architecture[37], originally designed for natural language processing, to image data. ViT utilizes image patch embedding, positional embeddings, transformer encoder, and classification token. While ViT achieved the state-of-the-art results in image classification when it appeared, it requires huge training datasets and it is computationally expensive for small datasets. DeiT has then successfully reduced the computational cost of training while maintaining accuracy by utilizing knowledge distillation [38] and data augmentation. We adopted DeiT due to our limited availability of computing resources. We make use of attention maps produced by DeiT that show the areas a machine learning model focuses on when it makes predictions, thereby improving the reliability of outputs. See the appendix A for further details about the Deit network we used.
@@ II.1 GW dataset generation
For the generation of CBC waveforms, we used PyCBC [39] with the SEOBNRv4_opt [40] waveform model. Our models were trained with parameters from Table 1. Although some parameters are restricted, we chose a simple configuration for the sake of simplicity, considering our main purpose.
Regarding the data generation procedure, we first created strain data by injecting a waveform into 4 second long data of stationary Gaussian noise colored by the LIGO design sensitivity (aLIGOZeroDetHighPower) [39]. During this process, the central time of the noise data was aligned with the merger time of the waveform. Then, we whitened the strain time series and cut the first and last 0.5 seconds of data to minimize possible edge effects. After the whitening process, we applied the constant-Q transform to create the spectrogram [41]. The constant-Q transform is a type of time-frequency transform similar to the Fourier transform. The constant-Q transform has high frequency resolution at low frequencies and high time resolution at high frequencies. Depending on the properties of data, the constant-Q transform is beneficial for catching detailed spectral structures than the usual Fourier transform.
@@
Table 1: Parameters used to create the training and validation datasets. $\mathcal{M}$, $q$, $\chi_{\text{eff}}$, $\theta_{JN}$, and $\psi$ are uniformly distributed. Source-frame component masses $m_{1}$$m_{2}$, are determined from $\mathcal{M}$ and $q$, while the dimensionless projections of the individual BH spins $\chi_{1}$ and $\chi_{2}$ are determined from $\chi_{\text{eff}}$. The luminosity distance $d_{L}$ and the sky position (right ascension $\alpha$ and declination $\delta$) are set to ensure a uniform distribution of gravitational wave sources within a sphere.
@@ II.2 Training and validation
As was mentioned in the introduction, we used Data-efficient Image Transformers (DeiT) [35] that is pre-trained on ImageNet [42] and we trained it on the dataset we created.
@@ II.3 Evaluation of uncertainty
To evaluate the uncertainty in the predictions of our models, we developed an uncertainty evaluation method based on a primitive Monte Carlo approach. See the appendix B for further explanation and possible use cases of our method. The flowchart of this method is shown in Fig. 3.
(1) Initially, a waveform was generated using parameters consistent with the estimates obtained from LALInference for GW150914. After injecting this waveform into stationary Gaussian noises colored by the same PSD as in Sec. II.1, a constant-Q transform was applied to create a spectrogram.
@@
(3) Subsequently, a waveform was generated based on the values estimated in (2). For parameters that were not estimated, the values were set to those used in (1). This waveform was injected into stationary Gaussian noises colored by the same PSD as in Sec. II.1, and a constant-Q transform was applied to create spectrograms. In this way, we created an uncertainty evaluation dataset that contains 1,000 spectrograms.
@@ III.1 Prediction results
Table 2 shows the 90% confidence intervals, while Fig. 4 and 5 illustrate the distributions of absolute errors. In Table 2, it can be seen that the predictions of our models were approximately consistent with those obtained by LALInference for GW150914. However, we did not analyze real GW events, and since the PSD used corresponds to the design sensitivity, this represents the result of a rough comparison. We also remark that since we used a frequentist approach while the results by LALInference are based on a Bayesian approach, two results are not necessarily expected to be the same.
@@
Figure 4: Histogram of absolute error for the effective spin estimation model. The absolute errors are calculated from the effective spins estimated in Sec. II.3-(4) and the estimated value in Sec. II.3-(2).
Figure 5: Histogram of absolute error for the chirp mass spin estimation model. The absolute errors are calculated from the chirp masses estimated in Sec. II.3-(4) and the estimated value in Sec. II.3-(2).
@@ III.2 Attention Map
We generated attention maps for the 1,000 spectrograms used in the uncertainty evaluation described in Sec. II.3. Each attention map we used is for the class token at the last self-attention layer of the network. To illustrate the overall trend, we created a composite figure by overlaying these attention maps (Fig. 6). It can be seen that the effective spin estimation model appears to focus primarily on the mid-frequency range, whereas the chirp mass estimation model focuses on the low-frequency range.
@@ IV.1 Method
We injected glitches into the simulated GW150914 data and applied the same uncertainty evaluation method as in Sec. II.3 to estimate the extent to which our model’s predictions would be biased if glitches were present in the simulated GW event. Furthermore, we examined the relationship between the Attention Map and the prediction results.
@@
$h_{\rm sg}=A_{\rm sg}e^{-\frac{(t-t_{0})^{2}}{\tau^{2}}}\cos\left(2\pi f_{0}(t-t_{0})\right)$ (1)
$A_{\rm sg}$ is an overall amplitude, $t_{0}$ and $f_{0}$ are the location of the sine-Gaussian in time and frequency, respectively, and $\tau$ is the damping time.
In real GW data analysis, it is problematic when glitches directly overlap GW signals as in the case of GW170817 [15]. However, injecting around the merger phase seemed likely to cause significant bias, thereby making it difficult to identify clear trends. Therefore, we selected $\tau=10^{-2}$ [s], $f_{0}=40$ [Hz] and $t_{0}=1.0$ [s] to configure the glitches to slightly overlap with the inspiral phase. For the amplitude $A_{\rm sg}$, we prepared three glitches with amplitudes of $0.5A$, $0.8A$, and $1A$, using the maximum amplitude $A$ of the simulated GW signal. In this manner, we generated spectrograms for three simulated GW events with injected glitches. Figure 8 presents an example.
@@ IV.2 Predictions
We compared the prediction results from Sec. III.1 with three simulated GW events with glitches of three different amplitudes.
Table 3 shows the 90% confidence intervals, while Fig. 9 and 10 illustrates the distributions of the predicted values. It can be seen that for both effective spin and chirp mass, glitches with larger amplitudes more strongly biased parameter estimation. This tendency is also commonly seen in conventional methods.
@@
$\chi_{\text{eff}}$ $09^{+0.07}_{-0.09} 0.0$$9^{+0.09}_{-0.07} 0.0$$7^{+0.08}_{-0.07} 0.0$$04^{+0.07}_{-0.05}$
$\mathcal{M}/M_{\odot}$ $27.6^{+0.8}_{-1.0}$ $30.5^{+0.8}_{-1.1}$ $2.6^{+0.6}_{-1.4} 33$$3.3^{+0.7}_{-1.3}$
@@ IV.3 Attention Map
Figures 11 and 12 shows the Attention Maps for a simulated GW150914 event without glitches and for three simulated GW events with varying glitch amplitudes. It can be seen that as the amplitude of the glitch increases, the attention around 1.0 seconds and 40 Hz, where the glitch was injected, also becomes larger.
@@
Furthermore, to quantify the impact of glitches to some extent, we calculated the total attention values for each time point. The distribution is shown in Figs. 13 and 14. The figures illustrate that as the amplitude of the glitch increases, the attention values around it also tend to increase.
@@ V Conclusion
In this paper, we proposed that attention maps enhance the reliability of the gravitational wave parameter estimation against transient noises in the detector output called glitches. Leveraging the power of the attention maps requires an inspection of each map for each event. If the detection rate of compact binary star mergers is small, as is the case currently, we can manually inspect the attention maps of each event. When the detection rate becomes larger, as is expected to happen in near future as the detectors’ sensitivities improve, we have to develop an automatic way of assessing the attention maps. One way to do that would be to check the consistency between the time of coalescence estimated from the matched filtering stage and the time when the attention map peaks.
As for other technical details, we used Gaussian noise to estimate parameter errors. We leave use of real data as future work. This work is intended to show that attention maps enhance the reliability of the machine learning technique in gravitational wave data analysis. For this purpose, we used a spectrogram as it is easier to understand than a time series or its Fourier transform. As a possible future direction, it would be interesting to use time series in addition to spectrogram to make a network learn phases of gravitational waves and glitches. Finally, we studied attention maps of the class token at the last output stage of the network. It would be interesting to use the attention roll-out that incorporates the information of the attention map of each layer [58].
Acknowledgements
We would like to thank Nobuyuki Kanda, Masaki Uematsu, Izumi Kaku, Takumi Fujimori, Ryuki Kawamoto, and Hiroyuki Nakano for helpful comments. We are grateful to Kota Tomita and Kyohei Takatani for their contribution in the early phase of this work. We would also like to thank Hirotaka Takahashi, Takashi Uchiyama, Yutaka Shikano, Shoichi Oshino and Yusuke Sakai for the valuable insights provided regarding the application of machine learning to gravitational wave data analysis. This work was supported by JSPS KAKENHI Grant No. 20H05639. We use PyTorch [59] for the implementation of our models. The plots are generated with matplotlib [60].
Appendix A Deit
In this work, we used the Data-efficient Image Transformers, specifically, the deit_small_distilled_patch16_224 model provided by the timm library, and we followed the default implementation for fine-tuning and inference. As described in Section II.B (Training and validation), we used DeiT models pre-trained on ImageNet. Therefore, we did not train a separate teacher network on the same tasks of mass and spin estimation. The basic architecture of the model is publicly available at the following site [61], and as shown there, in addition to the class token, the distillation token also contributes to the model’s predictions.
Appendix B Evaluation of uncertainties
In this appendix, we give the idea behind and possible use cases of our method of evaluating parameter uncertainties
We have used a simple frequentist approach. Let us explain it using a simple example. Suppose our data $x$ consists of a noise $n$ and signal $\mu$ like $x=n+\mu$. We may estimate the value of $\mu$ from $x$. Now, if we have lots of statistically independent $x$’s ($x_{1},x_{2},\cdots,x_{N}$) for a fixed $\mu$, then we obtain $N$ point-estimates of $\mu$ from which we obtain an error estimate of our point estimate of $\mu$. To realize this, we need to have lots of noise realizations $\{n_{i}\}_{i=1}^{N}$.
In the real-world application of our method, we propose that such noise realizations are obtained by injecting a simulated signal $h$ with fixed signal parameters $\hat{p}$ into time series starting at many different times, assuming that the statistical properties of the detector output time series are stable in time. Indeed, this is what is done to estimate the false alarm rate of gravitational wave events in the literature.
The question then is how we find $\hat{p}$. A possible flow of the data analysis would be the following. The first process is the matched filtering, with which one searches for gravitational wave events and, if detected, which gives us point estimates of the astrophysical parameters such as the masses of the stars. One usually uses, e.g., Markov-Chain Monte-Carlo or nested sampling techniques to estimate the parameter errors, but that takes huge computational costs. One of the problems we are tackling in this paper is this computational cost problem. As such, to some extent, we can assume that we are given point estimates of parameters from the search pipeline (matched filtering stage) before we start. We then use that set of estimated parameters to generate a simulated waveform and inject it into sets of time series at many different time stamps (where no signal is known to be present). We obtain astrophysical parameters from each time series and error estimates from them.